# A Review on Knowledge Graphs for Healthcare: Resources, Applications, and Promises

### v4

@@
Abstract.
Healthcare knowledge graphs (HKGs) are valuable tools for organizing biomedical concepts and their relationships. The recent advance of large language models (LLMs) has paved the way for building more comprehensive and accurate HKGs. This, in turn, can improve the reliability and evaluation of LLMs. However, the challenges and opportunities of HKGs are not fully understood, highlighting the need for detailed reviews. This work provides the first comprehensive review of HKGs, summarizing the pipeline and key techniques for HKG construction and successful HKG utilization in various health-related applications. Lastly, we highlight the opportunities for HKGs in the era of LLMs.
knowledge graph, healthcare, language models, multimodality, interpretability, trustworthy AI
$c$s: Applied computing Health informaticsccs: Applied computing Health care information systemsccs: Applied computing Bioinformaticsccs: Information systems Graph-based database modelsccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
Figure 1. The content overview of this review on healthcare knowledge graph.
@@ 1. Introduction
A knowledge graph (KG) is a data structure that captures the relationships between different entities and their attributes (Ji et al., 2021; Nicholson and Greene, 2020). KG models and integrates data from various sources, including structured and unstructured data, and has been studied to support a wide range of applications such as search engines (Wang et al., 2019b), recommendation systems (Wang et al., 2019a; Zhou et al., 2020), and question answering (Lin et al., 2019; Yasunaga et al., 2021; Yan et al., 2021; Kan et al., 2021). Healthcare Knowledge Graph (HKG) facilitates an interpretable representation of medical concepts, e.g., drugs and disease, as well as the relations among those medical concepts. This data structure enables the connection of contexts and enhances clinical research and decision-making (Santos et al., 2022; Chandak et al., 2023).
@@
On the data side, HKG is usually built on complex medical systems such as electronic health records, medical literature, clinical guidelines, and patient-generated data (Bouayad et al., 2017; Rajkomar et al., 2018). However, these data resources are often heterogeneous and distributed, which makes it challenging to integrate and analyze them effectively (Mehta and Pandit, 2018). This data heterogeneity can also lead to incomplete or inconsistent data representations, limiting their usefulness for downstream healthcare tasks (Dash et al., 2019). Additionally, the current use of domain-specific knowledge graphs may result in limited coverage and granularity of the knowledge captured across different levels. This hinders identifying correlations and relationships between medical concepts from multiple domains. These challenges highlight the need for continued research on HKGs to realize their full potential.
@@
On the modeling side, the construction of HKGs can be done either from scratch or by integrating existing dataset resources. Many crucial steps, such as entity and relation extraction, can be optimized with natural language processing tools and algorithms. Recently, there have been significant advancements in general domain knowledge extraction, thanks to pre-trained large language models (LLMs) such as BERT (Devlin et al., 2019), GPT Series (Brown et al., 2020), and others. These models revolutionize the field and make it possible to integrate heterogeneous medical data from various sources effectively. The use of pre-trained models has also led to the development of more accurate and comprehensive medical ontologies and taxonomies (Zhang et al., 2021d; Yu et al., 2020; Wang et al., 2021; Zeng et al., 2021; Xu et al., 2020). This allows for the evaluation of generated contents from LLMs and reduces LLM hallucination.
@@
A comprehensive healthcare knowledge graph has the potential to contribute to health research across various levels (Gyrard et al., 2018; Santos et al., 2022; Li et al., 2020). At the micro-scientific level, HKGs can help researchers identify new phenotypic and genotypic correlations and understand the underlying mechanisms of disease (Hassani-Pak and Rawlings, 2017), leading to more targeted and effective treatments (Seneviratne et al., 2021; Chandak et al., 2023). At the clinical care level, HKGs can be used to develop clinical decision support systems that provide clinicians with relevant information, improving clinical workflows and patient outcomes (Eberhardt et al., 2012; Castaneda et al., 2015). Therefore, a thorough review of existing literature on HKGs becomes an essential roadmap and invaluable resource to drive transformative advancements in the field.
This review is the first comprehensive overview of HKGs, covering contents shown in Figure 1. Specifically, in Section 2, we delve into the construction pipelines of HKGs, including both building from scratch and integration approaches, and highlight the key techniques employed in HKG construction. Additionally, we explore two standard utilization methods of HKGs, namely model-free and model-based approaches. In Section 3, we compile a comprehensive summary of existing HKG resources with their scopes and applications. Furthermore, Section 4 investigates the literature on mainstream health applications, offering an in-depth overview of the diverse use cases of HKGs in healthcare. Finally, we address promising research opportunities in the era of LLMs in Section 5.
2. Backgrounds
2.1. Knowledge Graphs for Healthcare (HKG)
Definition. A healthcare knowledge graph (HKG) is a domain-specific knowledge graph designed to capture medical concepts such as drugs, diseases, genes, phenotypes, and so on, and their relationships in a structured and semantic way.
Terminology. Our focus is on healthcare knowledge graphs (HKGs), which are structured, semantic representations of medical concepts and relationships. We also include ontologies and knowledge bases, which are commonly used in constructing HKGs. Ontologies define a set of concepts and categories in a domain, as well as the relationships between them, while knowledge bases store factual information about entities and their attributes. By covering these categories of terminology, we provide a comprehensive overview of the different types of resources available for organizing and representing medical knowledge in a structured and semantically rich manner.
@@ 2.1. Knowledge Graphs for Healthcare (HKG)
Figure 2. Detailed taxonomy of the background section of healthcare knowledge graphs.
2.2. HKG Construction
Healthcare knowledge graphs can be constructed from scratch or through the integration of existing data resources.
@@ 2.2. HKG Construction
Figure 3. The pipeline of constructing HKGs from scratch.
@@
  1. (1)
    The first step is to identify the scope and objectives. In most cases, researchers develop a schema (Guarino et al., 2009; Blagec et al., 2022) or use existing schemas (Guarino et al., 2009; Ashburner et al., 2000; Schriml et al., 2012; Bard et al., 2005) to serve as the formal and explicit specification of a domain, thus ensuring consistent, coherent, and aligned domain knowledge. Unlike the general domain KG, utilizing schemas is a common practice in HKG construction.
  2. (2)
@@
  3. (3)
    The third step is to extract and transform the data into a structured format. This step involves identifying medical entities and creating relationships between them via specialized biomedical Natural Language Processing (NLP) tools (Song et al., 2021; Xing et al., 2020; Hahn and Oleynik, 2020).
  4. (4)
    Next, researchers map the entities and relationships to the chosen ontologies with the help of thesauruses (Bodenreider, 2004) or terminologies (Donnelly et al., 2006; Hirsch et al., 2016). This step ensures that the knowledge graph is interoperable with other healthcare systems and facilitates data integration.
  5. (5)
    Until now, an initial KG has been built. The next step is to populate the KG to infer missing links between entities. This inference can be done using graph databases (Wang et al., 2020) or link prediction models (Bordes et al., 2013; Lu and Yang, 2022).
  6. (6)
@@
Constructing HKGs by Integration. Considering significant efforts have been paid to construct and curate HKGs from scratch, it is promising to integrate these data resources to avoid repetitive work. Healthcare KG integration (also called Healthcare KG fusion) refers to the processing of merging two or more HKGs into a single, more comprehensive graph (Himmelstein et al., 2017; Su et al., 2023; Youn et al., 2022). The integration process is challenging because different HKGs may use different terminologies, schemas, or data formats. To address these challenges, researchers have developed various techniques and algorithms for knowledge graph fusion, including ontology matching (Faria et al., 2014; He et al., 2022), schema alignment (Suchanek et al., 2011; Maaroufi et al., 2014), entity resolution (Bachman et al., 2018; Hu et al., 2021), and conflicts resolution (Ma et al., 2023). These methods aim to identify and reconcile the differences between KGs.
Techniques for HKG Constructions. Traditionally, each step of HKG construction involves one specially designated model. For instance, Hidden Markov Models and Recurrent Neural Networks are widely used for healthcare named entity recognition, relation extraction, and other sequence tagging tasks, while Translational Models and Graph Neural Networks are used for HKG completion and conflicts resolution tasks. Recently, large language models (LLMs) have shown great utility to serve as a uniform tool for constructing KGs (Ye et al., 2022). Several key steps of constructing KGs, such as named entity recognition (Liang et al., 2020; Chen et al., 2023; Huang et al., 2022; Liu et al., 2022b), relation extraction (Zhuang et al., 2022; Lu et al., 2022; Yang et al., 2022), entity linking (De Cao et al., [n. d.]; Mrini et al., 2022; Cho et al., 2022), and KG completion (geng2022disentangled; Saxena et al., 2022; Xie et al., 2022; Shen et al., 2022), have been successfully tackled by these large foundation models. Early explorations of construction HKG with large foundation models show that healthcare entity normalization (Zhang et al., 2023; Agrawal et al., 2022), healthcare entity recognition (Fries et al., 2021; Hu et al., 2023), healthcare entity linking (Zhu et al., 2022b), and healthcare knowledge fusion (Lu et al., 2023) can also be performed, without extensive training on expensive healthcare annotated corpus. On the other hand, researchers start to construct KGs under the open-world assumption (Shi and Weninger, 2018; Das et al., 2020; Niu et al., 2021; Li et al., 2022; Lu and Yang, 2022), thus getting rid of the dependency on pre-defined schemas and exhaustive entity&relation normalization. Although open-world KGs greatly increase the coverage, ensuring the quality of extracted knowledge is still an open research challenge, especially for explainable and trustworthy HKGs.
Table 1. Resource of existing healthcare knowledge graph (HKG).
@@
HetioNet (Himmelstein and Baranzini, 2015) 11 (e.g., drug, disease) 24 (e.g., drug-disease) #N: 47.0 K, #E: 2.3 M Medicinal Chemistry
DrKG (Ioannidis et al., 2020) 13 (e.g., disease, gene) 107 (e.g., disease-gene) #N: 97 K, #E: 5.8 M Medicinal Chemistry
PrimeKG (Chandak et al., 2023) 10 (e.g., phenotypes) 30 (e.g., disease-phenotype) #N: 129.4 K, #E: 8.1 M Medicinal Chemistry
@@
Gene Ontology [^1] http://geneontology.org/ (Ashburner et al., 2000) 3 (e.g., biological process) 4 (e.g., partOf) #N: 43 K, #E: 7544.6K Bioinformatics
KEGG [^2] https://www.genome.jp/kegg/ (Kanehisa and Goto, 2000) 16 (e.g., pathway) 4 (e.g., partOf) #N: 48.5 M, #E: unknown Bioinformatics
STRING [^3] https://string-db.org/ (Szklarczyk et al., 2023) 1 (e.g., protein) 4 (e.g., interactions) #N: 67.6 M, #E: 20 B Bioinformatics
Cell Ontology [^4] https://www.ebi.ac.uk/ols4/ontologies/cl (Diehl et al., 2016) 1 (i.e., cell type) 2 (e.g, subClassOf) #N: 2.7 K, #E: 15.9 K Bioinformatics
GEFA (Ranjan et al., 2022) 510 (e.g., kinases) 2 (e.g., drug-drug) #N: 0.5 K, #E: 30.1 K Drug Development
Reaction (Li and Chen, 2022) 2 (e.g., reactant & normal) 19 (e.g., reaction paths) #N: 2192.7 K, #E: 932.2 K Drug Development
ASICS (Jeong et al., 2022) 2 (e.g., reactant & product) 1 (e.g., reactions) #N: 1674.9 K, #E: 923.8 K Drug Development
@@
Hetionet (Jeong et al., 2022) 11 (e.g., biological process) 24 (e.g., disease–associates–gene) #N: 47.0 K, #E: 2250.2 K Drug Development
LBD-COVID (Zhang et al., 2021b) 1 (i.e., concept) 1 (i.e., SemMedDB relation) #N: 131.4 K, #E: 1016.1 K Drug Development
GP-KG (Gao et al., 2022) 7 (e.g., drug) 9 (e.g., disease–gene) #N: 61.1 K, #E: 1246.7 K Drug Development
DRKF (Zhang and Che, 2021) 4 (e.g., drug) 43 (e.g., drug-disease) #N: 12.5 K, #E: 165.9 K Drug Development
DDKG (Ghorbanali et al., 2023) 2 (i.e., drug & disease) 1 (e.g., drug-disease) #N: 551, #E: 2.7 K Drug Development
Disease Ontology [^5] https://disease-ontology.org/ (Schriml et al., 2012) 1 (i.e., disease) 2 (e.g., subClassOf) #N: 11.2 K, #E: 8.8 K Clinical Decision Support
DrugBank (Wishart et al., 2018) 4 (e.g., drug, pathway) 4 (e.g., drug-target) #N: 7.4 K, #E: 366.0 K Clinical Decision Support
KnowLife (Ernst et al., 2014) 6 (e.g., genes) 14 (e.g., gene-diseases) #N: 2.9 M, #E: 11.4 M Clinical Decision Support
PharmKG (Zheng et al., 2021) 3 (e.g., diseases) 3 (e.g., chemical-diseases) #N: 7601, #E: 500958 Clinical Decision Support
@@
ROBOKOP [^6] https://robokop.renci.org/ (Bizon et al., 2019) 54 (e.g., genes, drugs) 1064 (e.g., biolink, CHEBI) #N: 8.6M , #E: 130.4 M Clinical Decision Support
@@
iBKH [^7] https://github.com/wcm-wanglab/iBKH (Su et al., 2023) 11 (e.g., anatomy, disease) 18 (e.g., anatomy-gene) #N: 2.4 M, #E: 48.2 M Clinical Decision Support
2.3. HKG Utilization

### v5

@@
Abstract
Objective: This comprehensive review aims to provide an overview of the current state of Healthcare Knowledge Graphs (HKGs), including their construction, utilization models, and applications across various healthcare and biomedical research domains.
Methods: We thoroughly analyzed existing literature on HKGs, covering their construction methodologies, utilization techniques, and applications in basic science research, pharmaceutical research and development, clinical decision support, and public health. The review encompasses both model-free and model-based utilization approaches and the integration of HKGs with large language models (LLMs).
Results: We searched Google Scholar for relevant papers on HKGs and classified them into the following topics: HKG construction, HKG utilization, and their downstream applications in various domains. We also discussed their special challenges and the promise for future work.
Discussion: The review highlights the potential of HKGs to significantly impact biomedical research and clinical practice by integrating vast amounts of biomedical knowledge from multiple domains. The synergy between HKGs and LLMs offers promising opportunities for constructing more comprehensive knowledge graphs and improving the accuracy of healthcare applications.
Conclusions: HKGs have emerged as a powerful tool for structuring medical knowledge, with broad applications across biomedical research, clinical decision-making, and public health. This survey serves as a roadmap for future research and development in the field of HKGs, highlighting the potential of combining knowledge graphs with advanced machine learning models for healthcare transformation.
1 Introduction
@@
A knowledge graph (KG) is a data structure that captures the relationships between different entities and their attributes 1; 2. KG models and integrates data from various sources, including structured and unstructured data, and has been studied to support a wide range of applications such as search engines 3, recommendation systems 4; 5, and question answering 6; 7; 8; 9. Healthcare Knowledge Graph (HKG) facilitates an interpretable representation of medical concepts, e.g., drugs and disease, as well as the relations among those medical concepts. This data structure enables the connection of contexts and enhances clinical research and decision-making 10; 11.
@@
On the data side, HKG is usually built on complex medical systems such as electronic health records, medical literature, clinical guidelines, and patient-generated data 12; 13. However, these data resources are often heterogeneous and distributed, which makes it challenging to integrate and analyze them effectively 14. This data heterogeneity can also lead to incomplete or inconsistent data representations, limiting their usefulness for downstream healthcare tasks 15. Additionally, the current use of domain-specific KGs may result in limited coverage and granularity of the knowledge captured across different levels. This hinders identifying correlations and relationships between medical concepts from multiple domains. These challenges highlight the need for continued research on HKGs to realize their full potential.
@@
On the modeling side, the construction of HKGs can be done either from scratch or by integrating existing dataset resources. Many crucial steps, such as entity and relation extraction, can be done with NLP tools and algorithms. Recently, there have been significant advancements in general domain knowledge extraction, thanks to pre-trained large language models (LLMs) such as BERT 16, GPT Series 17, and others. These models revolutionize the field and make it possible to integrate heterogeneous medical data from various sources effectively. The use of pre-trained models has also led to the development of more accurate and comprehensive medical ontologies and taxonomies 18; 19; 20; 21; 22. This allows for the evaluation of generated contents from LLMs and reduces LLM hallucination.
@@
A comprehensive HKG has the potential to contribute to health research across various levels 23; 10; 24. At the micro-scientific level, HKGs can help researchers identify new phenotypic and genotypic correlations and understand the underlying mechanisms of disease 25, leading to more targeted and effective treatments 26; 11. At the clinical care level, HKGs can be used to develop clinical decision support systems that provide clinicians with relevant information, improving clinical workflows and patient outcomes 27; 28. Therefore, a thorough review of existing literature on HKGs becomes an essential roadmap and invaluable resource to drive transformative advancements in the field.
Statement of Significance: While prior reviews have focused on KGs tailored for specific healthcare tasks such as drug discovery 29, electronic health record (EHR) predictions 30, and ontologies 31, they often lack a holistic perspective across diverse healthcare domains. General surveys on KG construction 1 provide foundational insights but do not delve into the unique challenges and requirements of the healthcare context. Furthermore, some studies 32 concentrate solely on KG construction without exploring the critical dimension of knowledge utilization in healthcare applications. This work bridges these gaps by offering a comprehensive review that not only synthesizes knowledge graph construction techniques but also emphasizes their practical applications across various healthcare domains, providing a roadmap for future innovation.
Figure 1: The content overview of this review on healthcare knowledge graphs.
Literature Selection. We conducted a systematic literature review following the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) guidelines to identify and include studies relevant to the development and application of HKGs. Searches were conducted across multiple academic databases, including Google Scholar, PubMed, IEEE Xplore, ACM Digital Library, ACL Anthology, and Web of Science. Search terms were crafted to capture the scope of HKG research comprehensively. These terms included combinations of the following keywords using Boolean operators (AND, OR, NOT):
Healthcare Knowledge Graphs: “healthcare knowledge graph” “HKG” “Ontology” “medical knowledge graph”.
Applications: “clinical decision support” “drug repurposing” “pharmaceutical development” “clinical trial matching” “bioinformatics”.
Construction Techniques: “ontology integration” “entity resolution” “relation extraction” “knowledge graph embeddings”.
Integration with AI: “large language models” “machine learning” “natural language processing”.
Reference lists of key articles were manually searched to identify additional relevant studies. The search timeframe spanned from 2010 to 2024 with 362 articles in total. We then removed duplicate articles based on their titles and authors. After deduplication, we examined the remaining articles’ titles and abstracts for topic relevance. Finally, for each section, we conducted a detailed examination of the methodology described in each paper to ensure relevance, excluding those deemed irrelevant. After the filtering process, a total of 175 papers were retained.
Figure 2: The PRISMA diagram illustrating the literature search process.
2 Backgrounds
Healthcare Knowledge Graphs (HKGs). We focus on HKG, which is a structured representation of knowledge that captures entities (e.g., organizations, diseases, medications) and the relationships between them. KGs typically employ graph-based structures, where entities are represented as nodes and relationships as edges, and may use attributes or properties for additional contexts. These graphs can be enriched with contextual information and metadata, enabling complex relationships and data interconnectivity. Besides, we also include ontologies and knowledge bases, which are commonly used in constructing HKGs. An ontology is a formal model of a domain’s concepts, properties, and relationships, typically using a hierarchical or taxonomical structure, with an emphasis on semantic relation, and interoperability of different healthcare concepts. In contrast, KGs integrate data from heterogeneous sources, scale dynamically, and are well-suited for data-driven insights and probabilistic reasoning. By leveraging both ontologies and KGs, HKGs can support robust knowledge discovery and reasoning for healthcare applications. Furthermore, combining these resources with LLMs can enhance decision support by grounding LLM outputs in structured, contextualized knowledge and reduce the risk of hallucination. By covering these categories of terminology, we provide a comprehensive overview of the different types of resources available for organizing and representing medical knowledge in a structured and semantically rich manner.
@@ 2 Backgrounds
Figure 3: Detailed taxonomy of the background section of healthcare knowledge graphs.
2.1 HKG Construction
@@
Figure 4: The pipeline of constructing HKGs from scratch.
@@
  1. 1.
    The first step is to identify the scope and objectives. Researchers typically develop a schema 33; 34 or use existing schemas 33; 35; 36; 37 to clearly specify the domain, ensuring consistent and aligned knowledge. Unlike the general domain KG, utilizing schemas is a common practice in HKG construction.
  2. 2.
@@
  3. 3.
    The third step involves transforming the data into a structured format. This includes identifying medical entities and creating relationships between them via specialized biomedical NLP tools 38; 39; 40.
  4. 4.
    Entities and relationships are then mapped to selected ontologies using thesauruses 41 or terminologies 42; 43 to ensure the KG is compatible with other healthcare systems and supports data integration.
  5. 5.
    Once the initial KG is built, the next step is to fill in missing links between entities using graph databases 44 or link prediction models 45; 46.
  6. 6.
@@
Constructing HKGs by Integration. Considering significant efforts have been paid to construct and curate HKGs from scratch, it is promising to integrate these data resources to avoid repetitive work. HKG integration (also called HKG fusion) refers to the processing of merging two or more HKGs into a single, more comprehensive graph 47; 48; 49. The integration process is challenging because different HKGs may use different terminologies, schemas, or data formats. To address these challenges, researchers have developed various techniques and algorithms for KG fusion, including ontology matching 50; 51, schema alignment 52; 53, entity resolution 54; 55, and conflicts resolution 56. These methods aim to identify and reconcile the differences between KGs.
Techniques for HKG Constructions. Traditionally, each step of HKG construction involves one specially designated model. For instance, Hidden Markov Models and Recurrent Neural Networks are widely used for healthcare named entity recognition, relation extraction tasks, while Translational Models and Graph Neural Networks are used for HKG completion and conflict resolution tasks. Recently, LLMs have shown great utility to serve as a uniform tool for constructing KGs 57. Several key steps of constructing KGs, such as named entity recognition 58; 59; 60; 61, relation extraction 62; 63; 64, entity linking 65; 66; 67, and KG completion 68; 69; 70, have been successfully tackled by these large foundation models. Early explorations of construction HKG with large foundation models show that healthcare entity normalization 71; 72, healthcare entity recognition 73; 74, healthcare entity linking 75, and healthcare knowledge fusion 76 can also be performed, without extensive training on healthcare corpus. On the other hand, researchers start to construct KGs under the open-world assumption 77; 78; 79; 80; 46, thus getting rid of the dependency on pre-defined schemas and exhaustive entity&relation normalization.
Table 1: Resources of existing HKGs.
@@
HetioNet 81 11 (e.g., drug, disease) 24 (e.g., drug-disease) #N: 47.0 K, #E: 2.3 M Medicinal Chemistry
DrKG 82 13 (e.g., disease, gene) 107 (e.g., disease-gene) #N: 97 K, #E: 5.8 M Medicinal Chemistry
PrimeKG 11 10 (e.g., phenotypes) 30 (e.g., disease-phenotype) #N: 129.4 K, #E: 8.1 M Medicinal Chemistry
@@
Gene Ontology [^1] http://geneontology.org/ 35 3 (e.g., biological process) 4 (e.g., partOf) #N: 43 K, #E: 7544.6K Bioinformatics
KEGG [^2] https://www.genome.jp/kegg/ 83 16 (e.g., pathway) 4 (e.g., partOf) #N: 48.5 M, #E: unknown Bioinformatics
STRING [^3] https://string-db.org/ 84 1 (e.g., protein) 4 (e.g., interactions) #N: 67.6 M, #E: 20 B Bioinformatics
Cell Ontology [^4] https://www.ebi.ac.uk/ols4/ontologies/cl 85 1 (i.e., cell type) 2 (e.g, subClassOf) #N: 2.7 K, #E: 15.9 K Bioinformatics
GEFA 86 510 (e.g., kinases) 2 (e.g., drug-drug) #N: 0.5 K, #E: 30.1 K Drug Development
Reaction 87 2 (e.g., reactant & normal) 19 (e.g., reaction paths) #N: 2192.7 K, #E: 932.2 K Drug Development
ASICS 88 2 (e.g., reactant & product) 1 (e.g., reactions) #N: 1674.9 K, #E: 923.8 K Drug Development
@@
Hetionet 88 11 (e.g., biological process) 24 (e.g., disease–associates–gene) #N: 47.0 K, #E: 2250.2 K Drug Development
LBD-COVID 89 1 (i.e., concept) 1 (i.e., SemMedDB relation) #N: 131.4 K, #E: 1016.1 K Drug Development
GP-KG 90 7 (e.g., drug) 9 (e.g., disease–gene) #N: 61.1 K, #E: 1246.7 K Drug Development
DRKF 91 4 (e.g., drug) 43 (e.g., drug-disease) #N: 12.5 K, #E: 165.9 K Drug Development
DDKG 92 2 (i.e., drug & disease) 1 (e.g., drug-disease) #N: 551, #E: 2.7 K Drug Development
Disease Ontology [^5] https://disease-ontology.org/ 36 1 (i.e., disease) 2 (e.g., subClassOf) #N: 11.2 K, #E: 8.8 K Clinical Decision Support
DrugBank 93 4 (e.g., drug, pathway) 4 (e.g., drug-target) #N: 7.4 K, #E: 366.0 K Clinical Decision Support
KnowLife 94 6 (e.g., genes) 14 (e.g., gene-diseases) #N: 2.9 M, #E: 11.4 M Clinical Decision Support
PharmKG 95 3 (e.g., diseases) 3 (e.g., chemical-diseases) #N: 7601, #E: 500958 Clinical Decision Support
@@
ROBOKOP [^6] https://robokop.renci.org/ 96 54 (e.g., genes, drugs) 1064 (e.g., biolink, CHEBI) #N: 8.6M , #E: 130.4 M Clinical Decision Support
@@
iBKH [^7] https://github.com/wcm-wanglab/iBKH 48 11 (e.g., anatomy, disease) 18 (e.g., anatomy-gene) #N: 2.4 M, #E: 48.2 M Clinical Decision Support
2.2 HKG Utilization